OpenStudy (anonymous):

what hexadecimal number is equivalent to the binary number 111001110

OpenStudy (anonymous):

From last to beginning make the group of 4 \[ 111001110 = (0001)(1100)(1110)\] Now find the hexadecimal of these groups \[0001 = 1\] \[1100 = (1*2^3)+(1*2^2)+0+0= 8+4 = 12 = C\] \[1110 = (1* 2^3) + (1*2^2) + (1*2^1) + 0 = 8+4+2 = 14 = E\] Hexadecimal of \[ 111001110 = 1CE\]

OpenStudy (anonymous):

thanks but i don't get it i may have to go back and do some figuring.

OpenStudy (anonymous):

which part are you getting confused in?